ADNAN MUHAMMED

Grade 12,

if |z1|=1, |z2|=1,|1/z1 + 1/z2|=3/2 then find|z1+z2|?

if |z1|=1, |z2|=1,|1/z1 + 1/z2|=3/2 then find|z1+z2|?

Grade:12

1 Answers

Riddhish Bhalodia
8 years ago
This is quite simlpe
|1/z_1 + 1/z_2| = \frac{|z_1 +z_2|}{|z_1||z_2|} = |z_1 + z_2| = 3/2